OpenStudy (anonymous):

Subject: How does light tell us the speed of a distant object? Note: use the Doppler Shift for the following question/answer: 1. The sound produced by a passenger plucking a guitar string in a helicopter that is at rest relative to you has a wavelength of 3.50 m and a frequency of 100 Hz. If the helicopter is moving towards you at a velocity of 200 m/s, the frequency you hear is ___ Hz even though the passenger is still producing the same 100 Hz sound.

OpenStudy (anonymous):

\[FrequencyRecieved = \frac{ FrequencySent^{2} \times wavelength }{ FrequencySent \times wavelength + Velocity of Helicopter }\]

OpenStudy (anonymous):

I need it broken down with the answer because my math was not correct. I submitted it and it was incorrect. I know what the doppler shift formula is, but I flubbed the math part. Can anyone provide the answer?

OpenStudy (anonymous):

@JamesWolf can you maybe plug in the numbers so I can see which one I input in the wrong place?